# Increasing Profitability and Confidence by using Interpretable Model for Investment Decisions

### v1

@@ 1 Introduction
Decision-making is a complex cognitive process that involves analyzing various options and choosing the most optimal course of action. In the rapidly evolving landscape of financial markets, decision-making has taken on paramount significance, especially within the context of volatile markets such as the stock market. Investment decision in the stock market is a multifaceted process that requires navigation through the intricate web of fundamental analysis, market trends, economic indicators, and market sentiments [10]. Additionally, recognizing and capitalizing on opportunities promptly is essential for investors to leverage short-term market fluctuations in their investment strategies. Given the research dynamics of the stock market, such time-sensitive informed investment decisions towards buying and selling securities facilitate distinct market participants.
@@
In a conventional stock market environment, the equity research division fosters a collaborative and informative bond with market participants. Through regular research reports, trend forecasts, and personalized recommendations, the research team provides valuable insights into potential market opportunities. This, however, involves adopting various analytical methods and imperative forecasting to ensure that informed decisions are aligned with market participants’ investment goals [27]. When employing conventional research practices, the equity research division faces considerable obstacles while striving to publish research reports within a limited time frame following the closure of the daily stock market. The sheer volume of diverse market data is difficult to assess and synthesize effectively. It demands extensive cognitive effort and may involve the risk of biased decisions from research analysts of different backgrounds and experiences. Additionally, frequent price fluctuations in the market require timely investment recommendations for different groups of investors [8]. Each investor group has unique needs based on investment strategies, necessitating a careful and nuanced approach to provide relevant and actionable insights. For active traders, the analysis must highlight the real-time market trends and specific entry/exit trading points. Weekly traders, on the other hand, may seek more comprehensive analysis, including trends over a longer period with distinct preferences for certain technical indicators. With the integration of automation into traditional research processes, analysts are empowered to customize forecasted recommendations based on individual investor preferences, thereby elevating decision-making across different segments promptly.
@@
In recent years, the domain of financial forecasting has garnered considerable attention from the research community [22]. This interest can be attributed not only to the advancements in Artificial Intelligence (AI) but also to the abundant availability of financial data in digital form. Yet the role of AI is significant in the deployment of financial services in the industry like trading financial instruments [12], creditworthiness profiling [4], formulating credit or loan decisions [20], detecting fraud [26], predicting stock prices [17], assessing insurance claims [25], and managing debt collection [24]. The importance of utilizing an automated forecasting system that incorporates intelligent investment recommendations cannot be overstated. It serves as an indispensable tool for investors, relying on to make profitable decisions in the dynamic and complex financial market [21]. However, there are some serious impediments in the adoption of such end-to-end AI-based solutions that are entrusted with the power to make critical decisions, specifically in fields like health, defense, and finance where decision-making can affect people’s well-being. Regulatory authorities at the international level like the European Union have expressed concerns about the risks associated with the lack of ability to interpret and explain the decision-making process underlying these complex AI-based black-box models [1]. Given the highly regulated environments in which financial intermediaries, such as stock exchanges, operate, it has become crucial to consider the deployment of explainable modules when implementing solutions on a large scale. In time-critical scenarios, the equity research department’s responsibility towards diverse investors underscores the necessity for transparent and comprehensible methodologies, instilling confidence in those who rely on their reporting for investment decisions.
Lately, a novel concept of Explainable AI (XAI) has surfaced, which adds explainability and understandability to the outcomes of deployed automated models without compromising performance and accuracy. By far, data scientists are dealing with financial forecasting merely as an AI problem as highlighted by the Financial Stability Board [6], Germany’s Federal Financial Supervisory Authority [32], and policymakers in the European Union and Asia. Yet very few studies [11, 30, 31] have employed and addressed the importance of integrating transparent and trustworthy AI-based financial solutions with limited scope. Considering an existing research gap, this study centres on the exploration and application of a transparent and automated forecasting system, aiming to ensure adherence to regulatory standards within the financial sector.
The contribution of this study is multifold for addressing and incorporating interpretable decision-making in the context of investments in the stock exchange of Pakistan:
  1. [1.]
@@
    In a financial market, the identification of stakeholders and understanding their unique needs and rights is a crucial aspect of XAI. The proposed forecasting model considers and implements the variation among different types of investors.
@@
    In contrast to standard international stock exchange benchmarks such as S&P 500 and DJIA, the Pakistan Stock Exchange (PSX) operates within a distinct context. PSX’s performance is predominantly influenced by factors such as foreign investments, and political and economic stability, while key worldwide economic indicators such as consumer price index (CPI), unemployment ratio, etc. have a less significant impact on the volatility of Pakistan’s market. This highlights the need to consider the specific context and drivers of the PSX when analyzing its performance.
@@
    A proposed approach for end-users involves quantifying and modeling the factors that influence human economic decision-making, such as technical indicators. The goal of this study is to generate sound reports that can help users understand the rationale behind the model’s forecast.
The subsequent sections of the paper are structured in the following manner: Section 2 presents a review of notable research studies that have investigated the intersection of financial forecasting and explainability within the context of financial markets. This is followed by Section 3, which presents our proposed methodology. It integrates the explainability components in the solution addressing the requirements of different types of investors that can be addressed via equity research department. Section 4 provides the results and discussion as well as a case study that demonstrates the implementation of the proposed methodology. Finally, Section 5 concludes and highlights the paper findings along with the future direction of this research effort.
@@ 2 Literature Review
In recent years, several research studies on financial forecasting have been conducted, primarily aimed at easing the decision-making process for investors. This literature review places special emphasis on the efforts of researchers to provide explanations within the context of financial markets, specifically concerning the decision-making process resulting from proposed solutions. As opined by researchers [2, 3], complexity and unpredictability of financial markets make it challenging to accurately forecast future trends. Consequently, there has been an increasing interest in incorporating explainability into forecasting models to enhance transparency and interpretability for investors in the decision-making process.
@@
In another study, [6] presented an NLP-based approach for financial forecasting keeping in view the exogenous factors that affect the stock prices in a financial market. The goal was to achieve a methodology, both competitive in forecasting performance and explainable in the obtained output for the individual companies listed in the S&P 500 index. In the procedure, industry-specific lexicons were generated from the Dow Jones ‘Data, News, and Analytics (DNA) dataset. These lexicons were then refined through feature engineering, which involved selecting impactful words from the articles to capture certain statistical indicators associated with a company in a given time. The resulting list of features was used to train a decision tree classifier to predict whether a company’s stock price listed in the S&P 500 Index will rise or fall on the following day. Finally, the model explanation was achieved by presenting the rules to the users that are determined by the predictions drawn from the classifier. The explanation was complemented by a list of sentences, containing relevant lexicons words, selected from the groups of news articles considered in the feature extraction phase. The proposed model not only achieved a better F1 score in prediction but also may function as a superior tool for visualizing and explaining the results in the form of boolean logic conditions. However, their proposed solution falls short in incorporating technical indicators into the system to generate reports for end-users. To provide sound explanations, analysts must comprehend other technical insights beyond news, since mere words from news do not substantially alter the fluctuations in the stock market. Moreover, there is a need to cover the semantic relationship between the aggregated news over time for more accurate and comprehensive results.
@@
[32] investigated the production of post hoc plausible counterfactual explanations for an efficient financial solution using a robust transformer model. The objective of the study was to create an investment approach with significant returns by speculating on financial applications in the Fintech field. The authors of the study collected their dataset from Zephyr, a comprehensive database of real-world deal data, and were able to qualify over 4000 instances after preprocessing. Afterward, the authors utilized three methods - REP-SCD, RM-SCD, and INS-SCD to perform their prediction task using a classical transformer architecture on a given news article. As a next step, the explanation generated for the present prediction was flipped with appropriate words utilizing the three techniques that changed the prediction class of the original input. The final step involved generating counterfactual explanations using substitutions of grammatically plausible words. The results demonstrated superior accuracy and explanatory performance as compared to state-of-the-art techniques. The qualitative assessment of the more grammatically plausible explanations was evaluated by financial experts which received a score of 0.85. Although the authors focused on predicting mergers and acquisitions in financial companies with their proposed framework, however, they did not test it for canceled or pending deals. Additionally, relying solely on the positivity of actual news to determine the likelihood of a deal being completed may not adequately capture the complexities of real market dynamics [28].
@@
[30] in continuation to their prior work [19], incorporated explicit knowledge in the proposed sentiment-aware volatility forecasting (SAVING) model to forecast stock returns volatility. A knowledge-based approach that employed SenticNet was utilized in acquiring sentiment polarity from the social streaming on StockTwits for 10 US hotspot stocks. The proposed framework utilized the modern way of pricing derivatives such as the Black-Scholes model for stock volatility. A variety of models equipped with sentiment information from benchmark groups including linear and deep models were evaluated. Experiments showed the SAVING model outperformed many state-of-the-art volatility forecasting models including GARCH, VRNN, and NSVM. Even though the naïve generalized model has the expressive power of deep VRNNs and can capture bidirectional interaction between market sentiment and stock prices, the authors plan to enhance their approach by utilizing sentiment knowledge bases that are specific to the finance domain. Furthermore, the explainability module used in this study is limited to analyzing the polarity of individual sentences and requires a more sophisticated approach to be integrated for providing comprehensive explanations.
Stock market rumors refer to special information that begins to circulate in the stock market without confirmation. A new perspective of stock price linkage had been focused on by [9], highlighting the interplay between various types of investors in the stock market and how it can influence their risk preferences and behavior when it comes to rumors. To capture the complexity of the contagion process, the SCIR contagion mathematical model was constructed which incorporated the spread of risk from one investor to another, both within and across multiple networks of investors. By simulating the effects of rumors and certain probability mechanisms that can trigger risk contagion in the stock market on heterogeneous investors, the study aimed to assist regulators in devising effective strategies for mitigating financial risks. However, it is essential to conduct experiments with real market data to gain a better understanding of the emotional and psychological factors that influence investor behavior in actual trading scenarios.
@@
There are several key limitations in the existing research that need to be addressed, including but not limited to the insufficient adoption of popular technical indicators that are usually exercised during the technical or fundamental analysis for financial decision-making by the market stakeholders. Moreover, there is a lack in explicitly addressing varying stakeholders’ requirements while provisioning them with explanations. In a few studies, there is a gap in conducting experiments with real market data which is better for comprehending the investor’s behaviour and thus provisioning them with interpretations in a context. By addressing these gaps, the research community can establish a potential bridge between academic finance and the financial industry, enabling the development of more practical financial forecasting models, integrated with sophisticated explainability approaches. To the best of our knowledge, this is the first study of its kind that establishes the effort of providing visual explanations tailored to varying stock exchange investors. Moreover, the study integrates widely used technical indicators to analyze the data of Pakistan’s stock exchange, adding a new dimension to the research.
@@ 3 Methodology
In this research work, we acquired business-day financial data from the official website of the Pakistan Stock Exchange [23] as well as [14]. Our data spanned over a period of 10 years, specifically from 2012 to 2022, comprising approximately 2400 index entries, ensuring that we capture enough sample size to achieve our research objectives. The collected data encompasses various attributes of the KSE-100 Index, including the opening (O), highest (H), lowest (L), and closing (C) prices, as well as the traded volume (V) and the change in closing price $(\Delta C)$. These predictors provide the identification of trends and potential price reversals in the stock market which are crucial in making informed investment decisions. We can represent our raw feature space, $F_{d}$ as follows:
@@
To formulate the target feature, $Z_{t}$ we used a grid search technique to find optimal weights for each technical indicator-based strategy in S. A dictionary of weights W, is defined which contains the range of weights, denoted as $w_{i}fori\in\{1,2,3,4,5\}$ for each strategy. We then selected the set of weights with the highest accuracy as the overall optimal weights. A weighted voted function is then utilized with these optimal weights for each strategy to formulate a final trading signal for the investor for the following day.
@@
To validate our result, two investment strategies were designed, namely scalping and hold-and-wait to address the two types of investors i.e., daily active traders and weekly investors respectively. As for scalping, the accuracy of the one-day forecast is calculated by comparing the predicted target value with the actual price change for the following day from the curated stock data. The resulting target column, $Z_{t+1}\in[-3,3]$ is then used as a basis for making active trading investment decisions. Whereas, in a hold-and-wait investment strategy, the accuracy of predicted target values is assessed by comparing them with the actual price changes occurring over the course of the following week. The resulting target column, $Z_{t+5}\in[-3,3]$ is then used in the facilitation of short-term investment decisions by the investors. Algorithm 1. presents the pseudocode for the later investment strategy.
@@
We have compared our results with other state-of-the-art baseline approaches i.e., Random Forest and XGBoost to evaluate its performance and accuracy in forecasting the financial data of PSX, as well as its interpretability and explainability. The selection of these machine learning algorithms is motivated by the empirical findings, presented in the studies [13, 16], which demonstrate their effectiveness in producing valuable trading signals for portfolios characterized by frequent turnover and short holding periods over dual forecasting horizons. Given the dataset $F_{d}$ along with the technical indicators, feature engineering, followed by preprocessing of the data was performed in preparing the input for Random Forest Classifier. Heatmaps were utilized to identify and then eliminate the highly correlated features in the dataset. To avoid multicollinearity and redundant information, we removed $\Delta C_{d}$ i.e., the change in closing price each day which was overlapping with the $C_{d},$ the closing price. Furthermore, to avoid any feature domination, feature normalization was employed on the final dataset.

### v2

@@ 1 Introduction
Decision-making in financial markets is a complex cognitive process involving the analysis of various options to choose the most optimal course of action. In the rapidly evolving landscape of financial markets, this process has become increasingly significant, especially within the volatile environment of the stock market. Investment decisions in the stock market require navigating an intricate web of fundamental analysis, market trends, economic indicators, and market sentiments [10]. Additionally, recognizing and promptly capitalizing on opportunities is essential for investors to leverage short-term market fluctuations in their strategies. Given the dynamic nature of the stock market, time-sensitive, informed investment decisions towards buying and selling securities facilitate distinct market participants.
@@
In a conventional stock market environment, the equity research division fosters a collaborative and informative relationship with market participants. Through regular research reports, trend forecasts, and personalized recommendations, the research team provides valuable insights into potential market opportunities. This involves adopting various analytical methods and essential forecasting to ensure that informed decisions align with market participants’ investment goals [27]. However, the equity research division faces significant obstacles in publishing timely research reports due to the extensive cognitive effort required to assess and synthesize the sheer volume of diverse market data. This process can lead to biased decisions from analysts with different backgrounds and experiences. Additionally, frequent price fluctuations necessitate timely investment recommendations for different investor groups [8]. Each group has unique needs based on their investment strategies, requiring a nuanced approach to provide relevant and actionable insights. For active traders, the analysis must highlight real-time market trends and specific entry/exit trading points. Weekly traders may seek more comprehensive analysis, including longer-term trends and preferences for certain technical indicators. The integration of automation into traditional research processes empowers analysts to customize forecasted recommendations based on individual investor preferences, thereby enhancing decision-making across different segments promptly.
@@
In recent years, the domain of financial forecasting has garnered considerable attention from the research community [22]. This interest is driven by advancements in Artificial Intelligence (AI) and the abundant availability of financial data in digital form. AI plays a significant role in deploying financial services in various industries, including trading financial instruments [12], creditworthiness profiling [4], credit or loan decision-making [20], fraud detection [26], stock price prediction [17], insurance claim assessment [25], and debt collection management [24]. Utilizing an automated forecasting system with intelligent investment recommendations is indispensable for making profitable decisions in the dynamic and complex financial market [21]. However, the adoption of such AI-based solutions faces significant challenges, especially in fields like health, defence, and finance, where decision-making can impact well-being. Regulatory authorities, such as the European Union, have expressed concerns about the inability to interpret and explain the decision-making processes underlying complex AI-based black-box models [1]. Given the highly regulated environments in which financial intermediaries operate, it is crucial to deploy explainable modules for large-scale solutions. In time-critical scenarios, the equity research department’s responsibility towards diverse investors underscores the necessity for transparent and comprehensible methodologies, instilling confidence in those who rely on their reporting for investment decisions.
Recently, Explainable AI (XAI) has emerged, adding explainability and understandability to the outcomes of deployed automated models without compromising performance and accuracy. While financial forecasting has traditionally been treated as an AI problem [6], there are few studies [11, 30, 31] that have integrated transparent and trustworthy AI-based financial solutions. Addressing this research gap, this study focuses on exploring and applying a transparent and automated forecasting system to ensure adherence to regulatory standards within the financial sector.
Study Contributions
This study makes several key contributions to addressing and incorporating interpretable decision-making in the context of investments in the Pakistan Stock Exchange (PSX):
@@
    Stakeholder Identification and Needs Assessment: The proposed forecasting model considers and implements variations among different types of investors, addressing their unique needs and rights.
@@
    Context-Specific Analysis: Unlike international stock exchange benchmarks such as S&P 500 and DJIA, PSX operates within a distinct context influenced by factors like foreign investments, and political and economic stability, rather than global economic indicators like CPI and unemployment rates.
@@
    Quantifying Decision-Making Factors: The study involves quantifying and modeling the factors influencing human economic decision-making, using technical indicators to generate sound reports that help users understand the rationale behind the model’s forecasts.
The subsequent sections of the paper are structured in the following manner: Section 2 presents a review of notable research studies that have investigated the intersection of financial forecasting and explainability within the context of financial markets. This is followed by Section 3, which presents our proposed methodology. It integrates the explainability components in the solution addressing the requirements of different types of investors that can be addressed via the equity research department. Section 4 provides the results and discussion as well as a case study that demonstrates the implementation of the proposed methodology. Finally, Section 5 concludes and highlights the paper’s findings along with the future direction of this research effort.
@@ 2 Literature review
In recent years, several research studies on financial forecasting have been conducted, primarily aimed at easing the decision-making process for investors. This literature review places special emphasis on the efforts of researchers to provide explanations within the context of financial markets, specifically concerning the decision-making process resulting from proposed solutions. As opined by researchers [2, 3], the complexity and unpredictability of financial markets make it challenging to accurately forecast future trends. Consequently, there has been an increasing interest in incorporating explainability into forecasting models to enhance transparency and interpretability for investors in the decision-making process.
@@
In another study, [6] presented an NLP-based approach for financial forecasting keeping in view the exogenous factors that affect the stock prices in a financial market. The goal was to achieve a methodology, both competitive in forecasting performance and explainable in the obtained output for the individual companies listed in the S&P 500 index. In the procedure, industry-specific lexicons were generated from the Dow Jones ‘Data, News, and Analytics (DNA) dataset. These lexicons were then refined through feature engineering, which involved selecting impactful words from the articles to capture certain statistical indicators associated with a company in a given time. The resulting list of features was used to train a decision tree classifier to predict whether a company’s stock price listed in the S&P 500 Index will rise or fall on the following day. Finally, the model explanation was achieved by presenting the rules to the users that are determined by the predictions drawn from the classifier. The explanation was complemented by a list of sentences, containing relevant lexicon words, selected from the groups of news articles considered in the feature extraction phase. The proposed model not only achieved a better F1 score in prediction but also may function as a superior tool for visualizing and explaining the results in the form of boolean logic conditions. However, their proposed solution falls short in incorporating technical indicators into the system to generate reports for end-users. To provide sound explanations, analysts must comprehend other technical insights beyond news, since mere words from news do not substantially alter the fluctuations in the stock market. Moreover, there is a need to cover the semantic relationship between the aggregated news over time for more accurate and comprehensive results.
@@
[32] investigated the production of post hoc plausible counterfactual explanations for an efficient financial solution using a robust transformer model. The objective of the study was to create an investment approach with significant returns by speculating on financial applications in the Fintech field. The authors of the study collected their dataset from Zephyr, a comprehensive database of real-world deal data, and were able to qualify over 4000 instances after preprocessing. Afterwards, the authors utilized three methods - REP-SCD, RM-SCD, and INS-SCD to perform their prediction task using a classical transformer architecture on a given news article. As a next step, the explanation generated for the present prediction was flipped with appropriate words utilizing the three techniques that changed the prediction class of the original input. The final step involved generating counterfactual explanations using substitutions of grammatically plausible words. The results demonstrated superior accuracy and explanatory performance as compared to state-of-the-art techniques. The qualitative assessment of the more grammatically plausible explanations was evaluated by financial experts which received a score of 0.85. Although the authors focused on predicting mergers and acquisitions in financial companies with their proposed framework, however, they did not test it for cancelled or pending deals. Additionally, relying solely on the positivity of actual news to determine the likelihood of a deal being completed may not adequately capture the complexities of real market dynamics [28].
@@
[30] in continuation to their prior work [19], incorporated explicit knowledge in the proposed sentiment-aware volatility forecasting (SAVING) model to forecast stock returns volatility. A knowledge-based approach that employed SenticNet was utilized in acquiring sentiment polarity from the social streaming on StockTwits for 10 US hotspot stocks. The proposed framework utilized the modern way of pricing derivatives such as the Black-Scholes model for stock volatility. A variety of models equipped with sentiment information from benchmark groups including linear and deep models were evaluated. Experiments showed the SAVING model outperformed many state-of-the-art volatility forecasting models including GARCH, VRNN, and NSVM. Even though the naïve generalized model has the expressive power of deep VRNNs and can capture bidirectional interaction between market sentiment and stock prices, the authors plan to enhance their approach by utilizing sentiment knowledge bases that are specific to the finance domain. Furthermore, the explainability module used in this study is limited to analyzing the polarity of individual sentences and requires a more sophisticated approach to be integrated to provide comprehensive explanations.
Stock market rumors refer to special information that begins to circulate in the stock market without confirmation. A new perspective of stock price linkage has been focused on by [9], highlighting the interplay between various types of investors in the stock market and how it can influence their risk preferences and behavior when it comes to rumors. To capture the complexity of the contagion process, the SCIR contagion mathematical model was constructed which incorporated the spread of risk from one investor to another, both within and across multiple networks of investors. By simulating the effects of rumors and certain probability mechanisms that can trigger risk contagion in the stock market on heterogeneous investors, the study aimed to assist regulators in devising effective strategies for mitigating financial risks. However, it is essential to conduct experiments with real market data to gain a better understanding of the emotional and psychological factors that influence investor behavior in actual trading scenarios.
@@
There are several key limitations in the existing research that need to be addressed, including but not limited to the insufficient adoption of popular technical indicators that are usually exercised during the technical or fundamental analysis for financial decision-making by the market stakeholders. Moreover, there is a lack of explicitly addressing varying stakeholders’ requirements while providing them with explanations. In a few studies, there is a gap in conducting experiments with real market data which is better for comprehending the investor’s behaviour and thus provisioning them with interpretations in a context. By addressing these gaps, the research community can establish a potential bridge between academic finance and the financial industry, enabling the development of more practical financial forecasting models, integrated with sophisticated explainability approaches. To the best of our knowledge, this is the first study of its kind that establishes the effort of providing visual explanations tailored to varying stock exchange investors. Moreover, the study integrates widely used technical indicators to analyze the data of Pakistan’s stock exchange, adding a new dimension to the research.
@@ 3 Methodology
In this research work, we acquired business-day financial data from the official website of the Pakistan Stock Exchange [23] as well as [14]. Our data spanned over 10 years, specifically from 2012 to 2022, comprising approximately 2400 index entries, ensuring that we capture enough sample size to achieve our research objectives. The collected data encompasses various attributes of the KSE-100 Index, including the opening (O), highest (H), lowest (L), and closing (C) prices, as well as the traded volume (V) and the change in closing price $(\Delta C)$. These predictors provide the identification of trends and potential price reversals in the stock market which are crucial in making informed investment decisions. We can represent our raw feature space, $F_{d}$ as follows:
@@
To formulate the target feature, $Z_{t}$ we used a grid search technique to find optimal weights for each technical indicator-based strategy in S. A dictionary of weights W, is defined which contains the range of weights, denoted as $w_{i}for\textit{i}\in\{1,2,3,4,5\}$ for each strategy. We then selected the set of weights with the highest accuracy as the overall optimal weights. A weighted voted function is then utilized with these optimal weights for each strategy to formulate a final trading signal for the investor for the following day.
@@
To validate our result, two investment strategies were designed, namely scalping and hold-and-wait to address the two types of investors i.e., daily active traders and weekly investors respectively. As for scalping, the accuracy of the one-day forecast is calculated by comparing the predicted target value with the actual price change for the following day from the curated stock data. The resulting target column, $Z_{t+1}\in[-3,3]$ is then used as a basis for making active trading investment decisions. Whereas, in a hold-and-wait investment strategy, the accuracy of predicted target values is assessed by comparing them with the actual price changes occurring over the following week. The resulting target column, $Z_{t+5}\in[-3,3]$ is then used in the facilitation of short-term investment decisions by the investors. Algorithm 1. presents the pseudocode for the later investment strategy.
@@
  • $D\Leftarrow\{\text{ema55}_{d},\text{ema100}_{d},\text{ema200}_{d},\text{macd}_{d},\text{rsi}_{d}\}$ and $d=1,2,3……D$ where 'd'represents a specific day over a period 'D'
  • Strategy set S, representing trading signals
  • A dictionary of weights W
@@
  1. 1.
    Initialize a list of possible weight combinations, $W_{c}=S\times W=\{(s,w)\,|\,s\in S,w\in W\}$ is Cartesian product of two sets S and W.
  2. 2.
    Iterate over each set of weights in $W_{c}$, and do the following:
    • Calculate a weighted vote function, $wv(D,w_{cj})$ as in Eq. 8 for each row in D with $w_{cj},$ where $w_{cj}\in W_{c}$ to obtain $Z^{\prime}_{t+5}$ as intermediary target column formulation.
    • Call the $Agg\_signal\_5dayForecast$ function to calculate the accuracy for the weekly forecast using 20% of $Z^{\prime}_{t+5}$ and comparing with the next week’s change in closing price $\Delta C_{d+5}$ for $\Delta C_{d+5}\in F_{d}$ where 'd'represents a specific day within the time period.
  3. 3.
    Repeat step 2 for each weight combination $w_{cj}$ to get the best accuracy corresponding to the best weights $w_{best}$ in formulating optimized $Z_{t+5}$.
@@
We have compared our results with other state-of-the-art baseline approaches i.e., Random Forest and XGBoost to evaluate its performance and accuracy in forecasting the financial data of PSX, as well as its interpretability and explainability. The selection of these machine learning algorithms is motivated by the empirical findings, presented in the studies [13, 16], which demonstrate their effectiveness in producing valuable trading signals for portfolios characterized by frequent turnover and short holding periods over dual forecasting horizons. Given the dataset, $F_{d}$ along with the technical indicators, feature engineering, followed by preprocessing of the data was performed in preparing the input for Random Forest Classifier. Heatmaps were utilized to identify and then eliminate the highly correlated features in the dataset. To avoid multicollinearity and redundant information, we removed $\Delta C_{d}$ i.e., the change in closing price each day which was overlapping with the $C_{d},$ the closing price. Furthermore, to avoid any feature domination, feature normalization was employed on the final dataset.